OpenStudy (anonymous):

A map has a scale of 4 km to 3 inches. Which ratio of measures forms a proportion with the map's scale of km to inches? A:5 over 4. B:12 over 9. C:14 over 13. D:12 over 1.

OpenStudy (anonymous):

5/4 does not = 4km to 3in 14/13 does not = 4km to 3in 12/1 or 12 does not = 4km to 3in but ->*** 12/9 *does* = 4km to 3in

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank You So Much! <3

OpenStudy (anonymous):

Your Welcome!! :)